Find the slope of a line perpedicular to the line containing the points (-1, -11) and (9, 10)
Which line would be perpendicular to y = 3x + 2?
a) 3x-y=-1 b)y-3x=-1 c) x+3y=-1 d) 6x-2y=-1

PLEASE HELP ME!

Answer by nerdybill(7384)

Find the slope of a line perpedicular to the line containing the points (-1, -11) and (9, 10)
slope between two points:
(y2-y1)/(x2-x1)
(10-(-11))/(9-(-1))
(10+11)/(9+1)
21/10
a line perpendicular would be the negative reciprocal:
-10/21 (answer)
.
Which line would be perpendicular to y = 3x + 2?
a) 3x-y=-1 b)y-3x=-1 c) x+3y=-1 d) 6x-2y=-1
answer is (c)
because:
x+3y=-1
3y=-x-1
y=(-1/3)x-(1/3)
this equation has the negative reciprocal slope of 3